# C-LoRA: Contextual Low-Rank Adaptation for Uncertainty Estimation in Large Language Models

### v1

@@ 4.3 Robustness Under Distribution Shift
ACC $\uparrow$ MAP ${81.60{\pm_{0.40}}}$ ${73.88\pm_{0.27}}$ $67.67\pm_{1.52}$ ${41.00\pm_{3.60}}$ ${33.00\pm_{5.29}}$
MCD ${81.46{\pm_{0.92}}}$ ${75.00\pm_{2.13}}$ $68.69\pm_{0.85}$ ${39.00\pm_{3.00}}$ ${31.00\pm_{6.24}}$
Deep Ensemble ${82.20{\pm_{0.91}}}$ ${74.05\pm_{0.50}}$ $68.01\pm_{1.28}$ ${42.33\pm_{3.51}}$ ${27.66\pm_{2.51}}$
BLoB (M=0) ${81.85{\pm_{0.53}}}$ ${76.93\pm_{1.33}}$ $69.48\pm_{0.78}$ ${45.33\pm_{1.15}}$ ${26.66\pm_{4.61}}$
BLoB (M=10) ${81.44{\pm_{0.53}}}$ ${75.11\pm_{1.25}}$ $67.22\pm_{1.17}$ ${44.00\pm_{0.00}}$ ${33.66\pm_{2.08}}$
@@
C-LoRA (M=10) $78.26{\pm_{2.61}}$ ${74.29\pm_{2.90}}$ ${65.09\pm_{4.68}}$ $41.00\pm_{2.64}$ ${31.00\pm_{1.00}}$
ECE $\downarrow$ MAP ${15.30{\pm_{0.11}}}$ ${20.09\pm_{0.53}}$ $25.54\pm_{1.25}$ ${29.73\pm_{0.30}}$ ${36.22\pm_{3.60}}$
MCD ${14.34{\pm_{1.11}}}$ ${18.36\pm_{1.03}}$ $23.41\pm_{0.74}$ ${28.67\pm_{0.77}}$ ${36.53\pm_{4.30}}$
Deep Ensemble ${13.98{\pm_{1.12}}}$ ${16.89\pm_{0.80}}$ $20.90\pm_{1.05}$ ${16.10\pm_{2.22}}$ ${26.74\pm_{3.23}}$
BLoB (M=0) ${9.35{\pm_{1.08}}}$ ${11.70\pm_{0.62}}$ $15.76\pm_{1.42}$ ${14.12\pm_{4.05}}$ ${27.35\pm_{4.01}}$
BLoB (M=10) ${3.82{\pm_{0.96}}}$ ${\bf 5.74}\pm_{0.91}$ $\underline{9.77}\pm_{0.91}$ $\underline{12.63}\pm_{0.11}$ ${\bf 17.56}\pm_{2.81}$
@@
C-LoRA (M=10) $4.00{\pm_{0.84}}$ ${6.68\pm_{0.71}}$ ${\bf 8.83}\pm_{1.44}$ ${\bf 12.49}\pm_{1.18}$ $\underline{18.16}\pm_{1.52}$
NLL $\downarrow$ MAP ${1.04{\pm_{0.02}}}$ ${1.37\pm_{0.04}}$ $1.66\pm_{0.12}$ ${1.81\pm_{0.03}}$ ${1.86\pm_{0.04}}$
MCD ${1.01{\pm_{0.07}}}$ ${1.24\pm_{0.03}}$ $1.58\pm_{0.01}$ ${1.81\pm_{0.03}}$ ${1.88\pm_{0.10}}$
Deep Ensemble ${0.87{\pm_{0.03}}}$ ${0.94\pm_{0.06}}$ $1.15\pm_{0.09}$ ${1.45\pm_{0.01}}$ ${1.58\pm_{0.09}}$
BLoB (M=0) ${0.59{\pm_{0.02}}}$ ${0.72\pm_{0.01}}$ $0.95\pm_{0.05}$ ${1.41\pm_{0.02}}$ ${1.57\pm_{0.03}}$
BLoB (M=10) ${0.51{\pm_{0.01}}}$ ${\bf 0.63}\pm_{0.01}$ ${\bf 0.83}\pm_{0.02}$ $\underline{1.35}\pm_{0.01}$ $\underline{1.46}\pm_{0.01}$
@@ 4.4 Ablation Study: Impact of Contextual Module
C-LoRA (M=10) $4.00{\pm_{0.84}}$ ${\underline{6.68}\pm_{0.71}}$ ${\bf 8.83}\pm_{1.44}$ ${\bf 12.49}\pm_{1.18}$ ${\bf 18.16}\pm_{1.52}$
NLL $\downarrow$ FE (M=0) ${0.68{\pm_{0.04}}}$ ${0.87\pm_{0.01}}$ $1.03\pm_{0.01}$ ${1.53\pm_{0.05}}$ ${1.77\pm_{0.03}}$
FE (M=10) ${0.60{\pm_{0.03}}}$ ${0.78}\pm_{0.01}$ ${0.95}\pm_{0.02}$ ${1.46}\pm_{0.03}$ ${1.67}\pm_{0.02}$
@@
C-LoRA (M=10) $0.59{\pm_{0.05}}$ ${\underline{0.67}\pm_{0.02}}$ ${\underline{0.89}\pm_{0.09}}$ ${\bf 1.31}\pm_{0.02}$ ${\bf 1.42}\pm_{0.01}$

### v2

@@ 4.3 Robustness Under Distribution Shift
ACC $\uparrow$ MAP ${81.60{\pm_{0.40}}}$ $67.67\pm_{1.52}$ ${73.88\pm_{0.27}}$ ${41.00\pm_{3.60}}$ ${33.00\pm_{5.29}}$
MCD ${81.46{\pm_{0.92}}}$ $68.69\pm_{0.85}$ ${75.00\pm_{2.13}}$ ${39.00\pm_{3.00}}$ ${31.00\pm_{6.24}}$
Deep Ensemble ${82.20{\pm_{0.91}}}$ $68.01\pm_{1.28}$ ${74.05\pm_{0.50}}$ ${42.33\pm_{3.51}}$ ${27.66\pm_{2.51}}$
BLoB (M=0) ${81.85{\pm_{0.53}}}$ $69.48\pm_{0.78}$ ${76.93\pm_{1.33}}$ ${45.33\pm_{1.15}}$ ${26.66\pm_{4.61}}$
BLoB (M=10) ${81.44{\pm_{0.53}}}$ $67.22\pm_{1.17}$ ${75.11\pm_{1.25}}$ ${44.00\pm_{0.00}}$ ${33.66\pm_{2.08}}$
@@
C-LoRA (M=10) $78.26{\pm_{2.61}}$ ${65.09\pm_{4.68}}$ ${74.29\pm_{2.90}}$ $41.00\pm_{2.64}$ ${31.00\pm_{1.00}}$
ECE $\downarrow$ MAP ${15.30{\pm_{0.11}}}$ $25.54\pm_{1.25}$ ${20.09\pm_{0.53}}$ ${29.73\pm_{0.30}}$ ${36.22\pm_{3.60}}$
MCD ${14.34{\pm_{1.11}}}$ $23.41\pm_{0.74}$ ${18.36\pm_{1.03}}$ ${28.67\pm_{0.77}}$ ${36.53\pm_{4.30}}$
Deep Ensemble ${13.98{\pm_{1.12}}}$ $20.90\pm_{1.05}$ ${16.89\pm_{0.80}}$ ${16.10\pm_{2.22}}$ ${26.74\pm_{3.23}}$
BLoB (M=0) ${9.35{\pm_{1.08}}}$ $15.76\pm_{1.42}$ ${11.70\pm_{0.62}}$ ${14.12\pm_{4.05}}$ ${27.35\pm_{4.01}}$
BLoB (M=10) ${3.82{\pm_{0.96}}}$ $\underline{9.77}\pm_{0.91}$ ${\bf 5.74}\pm_{0.91}$ $\underline{12.63}\pm_{0.11}$ ${\bf 17.56}\pm_{2.81}$
@@
C-LoRA (M=10) $4.00{\pm_{0.84}}$ ${\bf 8.83}\pm_{1.44}$ ${6.68\pm_{0.71}}$ ${\bf 12.49}\pm_{1.18}$ $\underline{18.16}\pm_{1.52}$
NLL $\downarrow$ MAP ${1.04{\pm_{0.02}}}$ $1.66\pm_{0.12}$ ${1.37\pm_{0.04}}$ ${1.81\pm_{0.03}}$ ${1.86\pm_{0.04}}$
MCD ${1.01{\pm_{0.07}}}$ $1.58\pm_{0.01}$ ${1.24\pm_{0.03}}$ ${1.81\pm_{0.03}}$ ${1.88\pm_{0.10}}$
Deep Ensemble ${0.87{\pm_{0.03}}}$ $1.15\pm_{0.09}$ ${0.94\pm_{0.06}}$ ${1.45\pm_{0.01}}$ ${1.58\pm_{0.09}}$
BLoB (M=0) ${0.59{\pm_{0.02}}}$ $0.95\pm_{0.05}$ ${0.72\pm_{0.01}}$ ${1.41\pm_{0.02}}$ ${1.57\pm_{0.03}}$
BLoB (M=10) ${0.51{\pm_{0.01}}}$ ${\bf 0.83}\pm_{0.02}$ ${\bf 0.63}\pm_{0.01}$ $\underline{1.35}\pm_{0.01}$ $\underline{1.46}\pm_{0.01}$
@@ 4.4 Ablation Study: Impact of Contextual Module
C-LoRA (M=10) $4.00{\pm_{0.84}}$ ${\bf 8.83}\pm_{1.44}$ ${\underline{6.68}\pm_{0.71}}$ ${\bf 12.49}\pm_{1.18}$ ${\bf 18.16}\pm_{1.52}$
NLL $\downarrow$ FE (M=0) ${0.68{\pm_{0.04}}}$ $1.03\pm_{0.01}$ ${0.87\pm_{0.01}}$ ${1.53\pm_{0.05}}$ ${1.77\pm_{0.03}}$
FE (M=10) ${0.60{\pm_{0.03}}}$ ${0.95}\pm_{0.02}$ ${0.78}\pm_{0.01}$ ${1.46}\pm_{0.03}$ ${1.67}\pm_{0.02}$
@@
C-LoRA (M=10) $0.59{\pm_{0.05}}$ ${\underline{0.89}\pm_{0.09}}$ ${\underline{0.67}\pm_{0.02}}$ ${\bf 1.31}\pm_{0.02}$ ${\bf 1.42}\pm_{0.01}$